A 1000 kg sports car accelerates from 0 m/s to 30 m/s in 11.
What is the average power of the engine?
Answer #1
Concepts and reason

The concepts used in this problem are Newton’s equations of motion, Newton’s second law of motion, work done and the power.

First, calculate the displacement of the car using the Newton’s equation of motion. Then calculate the average power of the engine using the concept of work done and the power.

Fundamentals

Newton’s equations of motion:

The Newton’s equations of motion give the relation between initial speed, final speed, acceleration, distance and time of a motion. These equations describe the motion of an object.

Three equations of motion are:

vu=atv2u2=2ass=ut+12at2\begin{array}{l}\\v - u = at\\\\{v^2} - {u^2} = 2as\\\\s = ut + \frac{1}{2}a{t^2}\\\end{array}

Here, vvis the final speed, uuis the initial speed, aais the acceleration, ssis the distance and ttis the time.

Work done:

The work done in lifting an object from the ground is the product of force and displacement. The expression of work is:

W=FsW = Fs

Here, FFis the force andssis the displacement.

Power:

The energy per unit time is the power. The expression for power is:

P=EtP = \frac{E}{t}

Here, EEis the energy and ttis the time.

Newton’s Second Law of motion:

According to Newton’s second law of motion, the force on an object is equal to the product of mass and acceleration. It depends on mass and acceleration.

The expression of net force is:

Fnet=ma{F_{{\rm{net}}}} = ma

Here, mmis the mass and aais the acceleration.

The equations of motion used here are:

s=ut+12at2s = ut + \frac{1}{2}a{t^2} …… (1)

vu=atv - u = at …… (2)

The acceleration of the car is:

a=vuta = \frac{{v - u}}{t}

Substitute0m/s0{\rm{ m/s}}foruu, 30m/s30{\rm{ m/s}}forvv, 11s11{\rm{ s}}fortt in equation (3).

a=vut=30m/s0m/s11s=2.73m/s2\begin{array}{c}\\a = \frac{{v - u}}{t}\\\\ = \frac{{30{\rm{ m/s}} - 0{\rm{ m/s}}}}{{11{\rm{ s}}}}\\\\ = 2.73{\rm{ m/}}{{\rm{s}}^2}\\\end{array}

The displacement of the car is:

s=ut+12at2s = ut + \frac{1}{2}a{t^2}

Substitute0m/s0{\rm{ m/s}}foruu, 2.73m/s22.73{\rm{ m/}}{{\rm{s}}^2}foraa, 11s11{\rm{ s}}fortt in equation (3).

s=(0m/s)(11s)+12(2.73m/s2)(11s)2=165.2m\begin{array}{c}\\s = \left( {0{\rm{ m/s}}} \right)\left( {11{\rm{ s}}} \right) + \frac{1}{2}\left( {2.73{\rm{ m/}}{{\rm{s}}^2}} \right){\left( {11{\rm{ s}}} \right)^2}\\\\ = 165.2{\rm{ m}}\\\end{array}

The work done on the engine of the car is equal to the energy of the car according to work-energy theorem.

E=WE = W

E=FsE = Fs …… (3)

According to Newton’s second law of motion, the force is:

F=maF = ma …… (4)

Substitute equation (4) in equation (3).

E=masE = mas

Substitute1000kg{\rm{1000 kg}}formm, 2.73m/s2{\rm{2}}{\rm{.73 m/}}{{\rm{s}}^{\rm{2}}}forggand165.2m165.2{\rm{ m}}forss.

E=(1000kg)(2.73m/s2)(165.2m)=450996J\begin{array}{c}\\E = \left( {1000{\rm{ kg}}} \right)\left( {{\rm{2}}{\rm{.73 m/}}{{\rm{s}}^{\rm{2}}}} \right)\left( {165.2{\rm{ m}}} \right)\\\\ = 450996\;{\rm{J}}\\\end{array}

The power is:

P=EtP = \frac{E}{t}

Substitute450996J450996\;{\rm{J}}forEEand11s11\;{\rm{s}}fortt.

P=450996J11s=40999.6W=4.1×104W\begin{array}{c}\\P = \frac{{450996\;{\rm{J}}}}{{11\;{\rm{s}}}}\\\\ = 40999.6{\rm{ W}}\\\\ = {\bf{4}}{\bf{.1 \times 1}}{{\bf{0}}^{\bf{4}}}{\bf{ W}}\\\end{array}

Ans:

The average power of the engine is4.1×104W{\bf{4}}{\bf{.1 \times 1}}{{\bf{0}}^{\bf{4}}}{\bf{ W}}.